A shopkeeper sold a sofa set at 8 percent gain. Had he purchased it for 5 percent less and sold it for rs 36 less than what he actually sold it for,he would have gained 12.5 percent.For how much did the man purchase the sofa set?
Answer by mananth(16946)

A shopkeeper sold a sofa set at 8 percent gain. Had he purchased it for 5 percent less and sold it for rs 36 less than what he actually sold it for,he would have gained 12.5 percent.For how much did the man purchase the sofa set?
Let the purchase price of Sofa be Rs.x
He sold it for 8% gain.
He sold it for 1,08x
Cost at 5% less = 0.95x
Sold it for Rs36 less ie:1.08x-36
profit = (1.08x-36)-0.95x
= 0.13x -36
(0.13x-36)/0.95x = 0.125
0.13x -36 = 0.95x*0.125
0.13x-36=0.11875x
0.13x-0.11875x=36
solve for x
0.01125x =36
x = 3200
Cost price of the sofa set was Rs 3200




SP= 3200*1.08 =3456
5%less = 0.95*3200 =3040 CP
SP less 36 3456-36 = 3420
3420-3040 = 380
(380/3040)*100 =12.5%